# No Preference Left Behind: Group Distributional Preference Optimization

### v1

@@ Abstract
Preferences within a group of people are not uniform but follow a distribution. While existing alignment methods like Direct Preference Optimization (DPO) attempt to steer models to reflect human preferences, they struggle to capture the distributional pluralistic preferences within a group. These methods often skew toward dominant preferences, overlooking the diversity of opinions, especially when conflicting preferences arise. To address this issue, we propose Group Distribution Preference Optimization (GDPO), a novel framework that aligns language models with the distribution of preferences within a group by incorporating the concept of beliefs that shape individual preferences. GDPO calibrates a language model using statistical estimation of the group’s belief distribution and aligns the model with belief-conditioned preferences, offering a more inclusive alignment framework than traditional methods. In experiments using both synthetic controllable opinion generation and real-world movie review datasets, we show that DPO fails to align with the targeted belief distributions, while GDPO consistently reduces this alignment gap during training. Moreover, our evaluation metrics demonstrate that GDPO outperforms existing approaches in aligning with group distributional preferences, marking a significant advance in pluralistic alignment.  [^1] Our data and code will be released at https://github.com/BigBinnie/GDPO.
@@ 2 Related Work
Alignment techniques like RLHF and DPO face inherent limitations in achieving truly pluralistic AI, which aims to accommodate a broader range of human preferences (Sorensen et al., 2024). A promising approach to addressing this challenge involves aligning AI with multi-group preferences. Ramesh et al. (2024), Chakraborty et al. (2024), and Chen et al. (2024) propose frameworks for aligning LLMs with the preferences of different groups, acknowledging that populations can have divergent preferences. Meanwhile, other researchers focus on personalized alignment at the individual user level. Poddar et al. (2024), Shaikh et al. (2024), and Gao et al. (2024) explore methods that tailor LLM outputs based on users’ interaction histories and behaviors. Despite progress in pluralistic alignment, there is still limited research on how to represent and model diverse preferences within a single group so that multiple samplings of the model yield outputs as diverse as the population itself. Zhao et al. (2023) attempts to address this by aligning LLM outputs with group preference distributions using few-shot learning techniques, but their experiments are restricted to multiple-choice question-answer tasks, which fall short of capturing the complexity of real-world conversational language. Li et al. (2024) introduces distributional preference reward modeling to align with diverse preferences. However, it remains unclear whether the policy model can learn the same distribution through RLHF, and the training of RLHF has significant computational costs. To tackle these challenges, we propose GDPO , a reward-free alignment method, to align with distributional group preferences in controllable text generation settings.
@@ 4 Group Distributional Preference Optimization
Drawing inspiration from the language of thought hypothesis (LOTH) Fodor et al. (1975), which posits that thought occurs within a mental language, and formal epistemologists emphasize that beliefs exist on a spectrum, reflecting varying degrees of confidence or conviction Genin & Huber (2022), we define belief as the degree to which individuals agree with a particular stance. While individual preferences may vary depending on the context, beliefs can be represented as the extent of agreement with the statements in preference-related sentences. We discuss how we can conduct belief mining on preference datasets in Appendix §A.
@@
$\ell_{\text{pref}}(y_{c}\succ y_{r},b_{c},x)=$ $\log\sigma\left(\beta\log\frac{p_{\theta}(y_{c}\mid x,b_{c})}{p_{\text{ref}}(y_{c}\mid x,b_{c})}-\beta\log\frac{p_{\theta}(y_{r}\mid x,b_{c})}{p_{\text{ref}}(y_{r}\mid x,b_{c})}\right)$ (7)
@@ 5.1 Controllable Opinion Generation
We use synthetic data to simulate one-turn dialogues discussing global issues GlobalOpinionQA covers. Our synthetic data generation process consists of two steps: 1) dialogue data generation: for each multiple-choice question-answer pair in GlobalOpinionQA, we use GPT-3.5 to paraphrase the question in various styles. The answers are treated as distinct beliefs and corresponding responses are generated to express the opinions associated with the answer. The prompt for the data generation process is provided in Appendix C. Further, to improve belief calibration, we map the beliefs to six belief tokens representing varying degrees of agreement. The mapping table is provided in Appendix D; 2) conditional pairwise preference construction: we construct pairwise preference data based on country-specific statistics of opinions. The distribution of accepted beliefs aligns with the statistical distribution of opinions (i.e., answer distributions), while rejected beliefs are randomly sampled. Data examples are shown in Appendix E. To verify the performances on different distributions, we generate data from three different countries. They are the United States (US), Pakistan (PK), and S. Africa (SA). The dataset statistics are shown in Table 1. We prepare two different-sized datasets to train different-sized models.
@@
    Few-shot Prompts: We append a few examples to each prompt to show the preference distribution of the group. We test few-shot prompts both on our base models and GPT-4o. The prompt is shown in Appendix G.
@@ 5.1.1 Evaluation Metrics
  • JSD: Following the previous research work on GlobalOpinionQA Durmus et al. (2023); Zhao et al. (2023), we employ Jensen–Shannon Distance (JSD) to quantify the divergence between the ideal distribution and the learned distribution for each question. We report the average Jensen-Shannon distance over the test set ${\mathcal{D}}_{\text{test}}$, i.e., $\text{JSD}:=\frac{1}{N}\sum_{i=1}^{N}\textbf{Jensen–Shannon Distance}(p_{\theta}(b|x_{i}),p^{*}_{{\mathcal{B}},i})$.
@@ 5.1.2 Results and Analysis
We use movie reviews written by users from the Amazon Movie Review dataset [^2] https://snap.stanford.edu/data/web-Amazon.html to create the controllable movie review generation dataset. Our data is constructed in two steps: 1) prompt generation: for each movie, we retrieve background information about each movie from OMDB API [^3] https://www.omdbapi.com/, an open-source movie information database. Then we concatenate the movie’s metadata (e.g., title, genre, plot) with the movie title as the prompt; 2) conditional pairwise preference construction: for each movie review, we use the review as the accepted response and randomly sample one another review with a different rating for the same movie as the rejected response. Similar to controllable opinion generation, we construct two different-sized datasets. The statistics are shown in Table 4. Data examples are shown in Appendix F.
@@ 6 Conclusion and Limitations
Appendix A Discussion of No Belief Settings
@@ 7 Empathy and Understanding
For each data example in the dataset, we can annotate the belief by referencing these statements. As shown in Table 6, the preferences are all shaped by Statement 2.2. If the annotator chooses the chosen response, it indicates that the annotator’s belief aligns with the statement Do not promote illegal, harmful, or unethical activities. Conversely, if the annotator selects the rejected response, it suggests a disagreement with this belief. By aggregating all examples containing the same statement, we can estimate a belief distribution. Moreover, by regulating the belief distribution for safety-related topics, we can ensure that the generated responses adhere to safety standards, promoting only safe and ethical content.
@@
Appendix B Training Efficiency of GDPO
Appendix C Prompt of Data Synthesis
5. Don’t copy the exact words in the answer or use too uncommon words
The Question-Choices pair is:
Question: question
Choices: answers
please generate small talks which have [number_of_options].
Appendix D Class-Belief Mapping Table
Class tokens #Beliefs Belief Examples
B[0] 4 Not a moral issue, DK/Refused, Never heard of
Appendix E Data Examples of Controllable Opinion Generation
Example 1 Example 2
Belief distribution [0.7126, 0.2874] [0.26, 0.08, 0.26, 0.39, 0.01]
Accepted belief Yes Have not done, but might do
Rejected belief No Have done in the more distant past
Table 8: Controllable Opinion Generation Examples
Appendix F Data Examples of Controllable Movie Review Generation
Example 1 Example 2
Beliefs [1, 2, 3, 4, 5] [1, 2, 3, 4, 5]
Accept belief 5 1
Reject belief 4 4
Table 9: Controllable Review Generation Data Examples
Appendix G Few-shot Prompts Template
Question: Q1, Answer: <B1>R1
Question: Q2, Answer: <B2>R2
Question: Q3, Answer: <B3>R3
Question: Q4, Answer: <B4>R4
Question: Q5, Answer: <B5>R5
Answer:
Appendix H Evaluation by GPT-4o mini
BPC Evaluation of Controllable Opinion Generation.
Question: [question]
Answer: [answer]
Response in the following format:
Yes or No
One Sentence Explanation
BPC Evaluation of Controllable Review Generation.
1 means they disliked the film,
2 means they thought the film was okay or average,
3 means they loved the film.

### v2

@@ Abstract
Preferences within a group of people are not uniform but follow a distribution. While existing alignment methods like Direct Preference Optimization (DPO) attempt to steer models to reflect human preferences, they struggle to capture the distributional pluralistic preferences within a group. These methods often skew toward dominant preferences, overlooking the diversity of opinions, especially when conflicting preferences arise. To address this issue, we propose Group Distributional Preference Optimization (GDPO), a novel framework that aligns language models with the distribution of preferences within a group by incorporating the concept of beliefs that shape individual preferences. GDPO calibrates a language model using statistical estimation of the group’s belief distribution and aligns the model with belief-conditioned preferences, offering a more inclusive alignment framework than traditional methods. In experiments using both synthetic controllable opinion generation and real-world movie review datasets, we show that DPO fails to align with the targeted belief distributions, while GDPO consistently reduces this alignment gap during training. Moreover, our evaluation metrics demonstrate that GDPO outperforms existing approaches in aligning with group distributional preferences, marking a significant advance in pluralistic alignment. Our data and code are released at https://github.com/BigBinnie/GDPO.
@@ 2 Related Work
Alignment techniques like RLHF and DPO face inherent limitations in achieving truly pluralistic AI, which aims to accommodate a broader range of human preferences (Sorensen et al., 2024). A promising approach to addressing this challenge involves aligning AI with multi-group preferences. Ramesh et al. (2024), Chakraborty et al. (2024), and Chen et al. (2024) propose frameworks for aligning LLMs with the preferences of different groups, acknowledging that populations can have divergent preferences. Meanwhile, other researchers focus on personalized alignment at the individual user level. Poddar et al. (2024), Shaikh et al. (2024), Li et al. (2024b), and Gao et al. (2024) explore methods that tailor LLM outputs based on users’ identifiers, interaction histories, and behaviors. Despite progress in pluralistic alignment, there is still limited research on how to represent and model diverse preferences within a single group so that multiple samplings of the model yield outputs as diverse as the population itself. Zhao et al. (2023) attempts to address this by aligning LLM outputs with group preference distributions using few-shot learning techniques, but their experiments are restricted to multiple-choice question-answer tasks, which fall short of capturing the complexity of real-world conversational language. Li et al. (2024a) introduces distributional preference reward modeling to align with diverse preferences. However, it remains unclear whether the policy model can learn the same distribution through RLHF. Siththaranjan et al. (2023) proposes applying the distributional preference learning to RLHF. However, RLHF has significant computational costs. To tackle these challenges, we propose GDPO, a reward-free alignment method, to align with distributional group preferences in controllable generation settings.
@@ 4 Group Distributional Preference Optimization
Drawing inspiration from the language of thought hypothesis (LOTH) Fodor et al. (1975), which posits that thought occurs within a mental language, and formal epistemologists emphasize that beliefs exist on a spectrum, reflecting varying degrees of confidence or conviction Genin & Huber (2022), we define belief as the degree to which individuals agree with a particular stance. While individual preferences may vary depending on the context, beliefs can be represented as the extent of agreement with the statements in preference-related sentences. We discuss how we can conduct belief mining on preference datasets in Appendix §I.
@@
$\ell_{\text{pref}}(y_{c}\succ y_{r},b_{c},x)=-$ $\log\sigma\left(\beta\log\frac{p_{\theta}(y_{c}\mid x,b_{c})}{p_{\text{ref}}(y_{c}\mid x,b_{c})}-\beta\log\frac{p_{\theta}(y_{r}\mid x,b_{c})}{p_{\text{ref}}(y_{r}\mid x,b_{c})}\right)$ (7)
@@ 5.1 Controllable Opinion Generation
We use synthetic data to simulate one-turn dialogues discussing global issues GlobalOpinionQA covers. Our synthetic data generation process consists of two steps: 1) dialogue data generation: for each multiple-choice question-answer pair in GlobalOpinionQA, we use GPT-3.5 to paraphrase the question in various styles. The answers are treated as distinct beliefs and corresponding responses are generated to express the opinions associated with the answer. The prompt for the data generation process is provided in Appendix A. Further, to improve belief calibration, we map the beliefs to six belief tokens representing varying degrees of agreement. The mapping table is provided in Appendix B; 2) conditional pairwise preference construction: we construct pairwise preference data based on country-specific statistics of opinions. The distribution of accepted beliefs aligns with the statistical distribution of opinions (i.e., answer distributions), while rejected beliefs are randomly sampled. Data examples are shown in Appendix C. To verify the performances on different distributions, we generate data from three different countries. They are the United States (US), Pakistan (PK), and S. Africa (SA). The dataset statistics are shown in Table 1. We prepare two different-sized datasets to train different-sized models.
@@
    Few-shot Prompts: We append a few examples to each prompt to show the preference distribution of the group. We test few-shot prompts both on our base models and GPT-4o. The prompt is shown in Appendix E.
@@ 5.1.1 Evaluation Metrics
  • JSD: Following the previous research work on GlobalOpinionQA Durmus et al. (2023); Zhao et al. (2023), we employ Jensen–Shannon Distance (JSD) to quantify the divergence between the ideal distribution and the learned distribution for each question. We report the average Jensen-Shannon distance over the test set ${\mathcal{D}}_{\text{test}}$, i.e.,
    $\text{JSD}:=\frac{1}{N}\sum_{i=1}^{N}\textbf{JSD}(p_{\theta}(b|x_{i}),p^{*}_{{\mathcal{B}},i})$.
@@ 5.1.2 Results and Analysis
We use movie reviews written by users from the Amazon Movie Review dataset [^1] https://snap.stanford.edu/data/web-Amazon.html to create the controllable movie review generation dataset. Our data is constructed in two steps: 1) prompt generation: for each movie, we retrieve background information about each movie from OMDB API [^2] https://www.omdbapi.com/, an open-source movie information database. Then we concatenate the movie’s metadata (e.g., title, genre, plot) with the movie title as the prompt; 2) conditional pairwise preference construction: for each movie review, we use the review as the accepted response and randomly sample one another review with a different rating for the same movie as the rejected response. Similar to controllable opinion generation, we construct two different-sized datasets. The statistics are shown in Table 4. Data examples are shown in Appendix D.
@@ 6 Conclusion and Limitations
Acknowledgments
Yao and Hu are supported by the American Family Insurance Data Science Award and the Wisconsin Alumni Research Foundation. Jiang is partially supported by the National Science Foundation (IIS-2438420). This research is partially supported by the NVIDIA Academic Grant Program and the Microsoft Accelerating Foundation Models Research Program. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Science Foundation.
Appendix A Prompt of Data Synthesis
5. Don’t copy the exact words in the answer or use too uncommon words
The Question-Choices pair is:
Question: question
Choices: answers
please generate small talks which have [number_of_options].
Appendix B Class-Belief Mapping Table
Class tokens #Beliefs Belief Examples
B[0] 4 Not a moral issue, DK/Refused, Never heard of
Appendix C Data Examples of Controllable Opinion Generation
Example 1 Example 2
Belief distribution [0.7126, 0.2874] [0.26, 0.08, 0.26, 0.39, 0.01]
Accepted belief Yes Have not done, but might do
Rejected belief No Have done in the more distant past
Table 6: Controllable Opinion Generation Examples
Appendix D Data Examples of Controllable Movie Review Generation
Example 1 Example 2
Beliefs [1, 2, 3, 4, 5] [1, 2, 3, 4, 5]
Accept belief 5 1
Reject belief 4 4
Table 7: Controllable Review Generation Data Examples
Appendix E Few-shot Prompts Template
Question: Q1, Answer: <B1>R1
Question: Q2, Answer: <B2>R2
Question: Q3, Answer: <B3>R3
Question: Q4, Answer: <B4>R4
Question: Q5, Answer: <B5>R5
Answer:
Appendix F Training Efficiency of GDPO
Appendix G Experiment Setup.
Appendix H Evaluation by GPT-4o mini
BPC Evaluation of Controllable Opinion Generation.
Question: [question]
Answer: [answer]
Response in the following format:
Yes or No
One Sentence Explanation
BPC Evaluation of Controllable Review Generation.
1 means they disliked the film,
2 means they thought the film was okay or average,
3 means they loved the film.
Appendix I Discussion of No Belief Settings
@@ 7 Empathy and Understanding
For each data example in the dataset, we can annotate the belief by referencing these statements. As shown in Table 6, the preferences are all shaped by Statement 2.2. If the annotator chooses the chosen response, it indicates that the annotator’s belief aligns with the statement Do not promote illegal, harmful, or unethical activities. Conversely, if the annotator selects the rejected response, it suggests a disagreement with this belief. By aggregating all examples containing the same statement, we can estimate a belief distribution. Moreover, by regulating the belief distribution for safety-related topics, we can ensure that the generated responses adhere to safety standards, promoting only safe and ethical content. Further exploration of improved methods for modeling belief in latent variables is both intriguing and promising.
@@
Appendix J Generalization to KTO
In this section, we demonstrate how GDPO could be applied to other alignment methods, such as KTO (Ethayarajh et al. (2024)). KTO loss is Eq. 8, which calculates the losses of desirable and undesirable samples separately.
$L_{\mathrm{KTO}}\left(\pi_{\theta},\pi_{\mathrm{ref}}\right)=\mathbb{E}_{x,y\sim D}\left[\lambda_{y}-v(x,y)\right]$ (8)
where
$r_{\theta}(x,y)=\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}$ (9)
$z_{0}=\operatorname{KL}\left(\pi_{\theta}(y^{\prime}\mid x)\,\|\,\pi_{\mathrm{ref}}(y^{\prime}\mid x)\right)\$ (10)
$v(x,y)=\begin{cases}\lambda_{D}\sigma\left(\beta\left(r_{\theta}(x,y)-z_{0}\right)\right),&\text{if }y\sim y_{\text{desirable}}\mid x\\ \lambda_{U}\sigma\left(\beta\left(z_{0}-r_{\theta}(x,y)\right)\right),&\text{if }y\sim y_{\text{undesirable}}\mid x\end{cases}$ (11)
When applying GDPO to KTO, we use KTO loss as the belief-conditioned preference alignment loss by changing the reward in Eq. 9 to be conditioned on belief. The $v(x,y)$ is shown in Eq. 12.
$v(x,y)=\begin{cases}\lambda_{D}\sigma\left(\beta\left(r_{\theta}(x,b_{\text{desirable}},y)-z_{0}\right)\right),&\text{if }y\sim y_{\text{desirable}}\mid x\\ \lambda_{U}\sigma\left(\beta\left(z_{0}-r_{\theta}(x,b_{\text{desirable}},y)\right)\right),&\text{if }y\sim y_{\text{undesirable}}\mid x\end{cases}$ (12)